OpenStudy (anonymous):

help me !!! 2xy+3x+4x+7xy=90 find x in terms of y

OpenStudy (lalaly):

you find them

OpenStudy (anonymous):

HOW ???

OpenStudy (mimi_x3):

unfathomable :P

OpenStudy (anonymous):

mememememememe will help

OpenStudy (lalaly):

=D

OpenStudy (anonymous):

virus plz help me

OpenStudy (mimi_x3):

The virus has detected that the equation entered is unfathomable :P

OpenStudy (mr.math):

Take x out as a common factor.

OpenStudy (anonymous):

@mr mats ... how ... i m new to algebra

OpenStudy (mr.math):

\[2xy+3x+4x+7xy=x(2y+3+4+7y)=x(9y+7).\]

OpenStudy (anonymous):

2xy+3x+4x+7xy=90 x(2y+3+4+7y)=90 x=90/9y+7

OpenStudy (mr.math):

Now dived both sides of your equation by 9y+7, as ishaan did.

OpenStudy (angela210793):